# SynCoGen: Synthesizable 3D Molecule Generation via Joint Reaction and Coordinate Modeling

### v1

@@ A.1 Building Blocks and Reactions
A.2 Graph Generation
Helper definitions.
We annotate each building block with its reaction center atom indices $\mathcal{V}(b)\subseteq V(b)$ and its and each intrinsic atom-level graph by $H(b)\;:=\;\bigl{(}V(b),\,L(b)\bigr{)}$, where $V(b)$ is the set of atoms in $b$ and $L(b)\subseteq V(b)\times V(b)$ is the set of covalent bonds internal to the block. Each reaction template $r$ is annotated with a Boolean tuple $\bigl{(}(l_{A}(r),\,l_{B}(r)\,\bigr{)}\in\{0,1\}^{2}$ describing whether reagent $A$ or reagent $B$ in $r$, respectively, contains a leaving atom.
Given the current atom graph $G_{a}=(V_{a},L_{a})$ and an atom $v\in V_{a}$ of degree 1, the routine $\textsc{UniqueNeighbor}(v)$ returns the single atom $u\in V_{a}$ such that $(u,v)\in L_{a}$. Throughout the vocabulary, every leaving-group center has exactly one neighbour.
A reaction template $r$ is considered compatible with $(b_{i},v)$ and $(\tilde{b},\tilde{v})$ if it queries for first and second reagent substructures that match $(b_{i},v)$ and $(\tilde{b},\tilde{v})$, respectively.
Figure 5: List of building blocks, their respective reaction centers (in red), and their canonical SMILES representation.
Figure 6: List of chemical reactions used to connect building blocks and their SMARTS representation. Newly formed bonds are highlighted in pink.
Algorithm 3 Fragment-by-fragment assembly with Couple
Inputs: vocab $\mathcal{B}$, reactions $\mathcal{R}$, depth limit $T$
Output: atom graph $G_{a}$, building block graph $G_{f}=(X,E)$
1:function Couple($G_{a},\;b_{i},\;\tilde{b},\;r,\;(v_{i},\tilde{v})$)
2:     append all atoms and bonds of $H(\tilde{b})$ to $G_{a}$triangleright 1. Handle leaving groups
3:     if $l_{A}(r)=1$ then $\triangleright$ $v_{i}$ leaves in reagent A
4:         $u_{i}\leftarrow$ UniqueNeighbor($v_{i}$)
5:         delete atom $v_{i}$ (and its bond) from $G_{a}$
6:         $v_{i}\leftarrow u_{i}$triangleright reroute to neighbour
7:     end if
8:     if $l_{B}(r)=1$ then $\triangleright$ $\tilde{v}$ leaves in reagent B
9:         $u_{\text{t}}\leftarrow$ UniqueNeighbor($\tilde{v}$)
10:         delete atom $\tilde{v}$ (and its bond) from $G_{a}$
11:         $\tilde{v}\leftarrow u_{\text{t}}$triangleright reroute to neighbour
12:     end if
13:     add covalent bond between $v_{i}$ and $\tilde{v}$triangleright 2. Add the cross-bond
14:     return $G_{a}$
15:end function
16:$b_{0}\leftarrow\text{UniformPick}(\mathcal{B})$; $G_{a}\leftarrow H(b_{0})$; $G_{f}\leftarrow(b_{0})$
17:for $t=1$ to $T$ do
18:     $L\leftarrow$ enumerate compatible 5-tuples $\langle b_{i},v,r,\tilde{b},\tilde{v}\rangle$
19:     if $L=\varnothing$ then break
20:     end if
21:     $(b_{i},v,r,\tilde{b},\tilde{v})\leftarrow\text{UniformPick}(L)$
22:     $e\leftarrow(r,v,\tilde{v})$
23:     $G_{a}\leftarrow\textsc{Couple}(G_{a},\,b_{i},\,\tilde{b},\,r,(v,\tilde{v}))$
@@ A.2 Graph Generation
25:end for
26:return $(G_{a},G_{f})$
Appendix B Method Details
B.1 Building Block-Level Representations
Let $X\in\{0,1\}^{N\times|\mathcal{B}|+1}$ be a one-hot matrix where the $i^{\text{th}}$ row encodes the identity of the $i^{\text{th}}$ building block, and let $E\;\in\;\{0,1\}^{N\times N\times|\mathcal{R}|V_{max}^{2}+2},$ where $V_{max}=\max_{b}|\mathcal{V}(b)|$. A non-zero entry $E_{ijr(v_{i},v_{j})}=1$ signals that block $i$ (center $v_{i}$) couples to block $j$ (center $v_{j}$) via reaction $r$. Graphs $(X,E)$ belonging to molecules containing $n<N$ building blocks are padded to $N$.
Reserved Channels.
where $\pi_{X}$ (resp. $\pi_{E}$) is the one-hot vector whose single 1-entry corresponds to the masked node (resp. edge) channel. Besides the masked channel, we keep a dedicated no-edge channel, encoded by the one-hot vector
$\lambda_{E}\in\{0,1\}^{|\mathcal{R}|\,V_{\max}^{2}},$ (8)
so every edge slot may take one of three mutually exclusive states: a concrete coupling label, the no-edge token $\lambda_{E}$, or the masked token $\pi_{E}$.
B.2 Atom-Level Representations
The SemlaFlow [Irwin et al., 2025] architecture propagates and updates invariant and equivariant features at the atom level. To ensure consistency with this framework, we calculate for each input graph ($X_{t},E_{t}$) atom-level one-hot atom and bond features. Crucially, these features must be flexible to arbitrary masking present in $X_{t}$ and $E_{t}$. With this in mind we set each atom feature $X^{atom}_{t}[i,a]$ to a concatenation of one-hot encodings
[…]
where $\delta_{\mathrm{sym}}(i,a)$ is the one-hot vector over possible atom types (C, N, O, B, F, Cl, Br, S, [MASK]) and ring$(i,a)$ denotes whether or not the atom is a member of a ring. Similarly, we calculate a bond feature matrix
B.3 Training Algorithm
Algorithm 4 Training step for SynCoGen
$t\sim\mathcal{U}(0,1)$
2:$(X_{t},E_{t})\leftarrow q_{t}(X_{0},E_{0})$
3:$C_{1}\sim\mathcal{N}(0,I)$
4:$(\tilde{C}_{0},\tilde{C}_{t})\leftarrow\textsc{Pair}(C_{0},S_{0},C_{1},t,X_{t})$triangleright center and interpolate coordinates (Algorithm 1)
6:$\mathcal{L}\leftarrow\mathcal{L}_{\text{graph}}+\mathcal{L}_{\text{MSE}}+\mathcal{L}_{\text{pair}}$triangleright total loss (Section B.11)
7:$\theta\leftarrow\theta-\eta\,\nabla_{\theta}\mathcal{L}$
B.4 Sampling Algorithm
Algorithm 5 Sampling procedure for SynCoGen
1:$n\!\sim\!\text{Cat}(\pi_{\text{frag}})$;  $(X_{1},E_{1})\!\leftarrow(\pi_{X},\pi_{E})$;  $S_{1}[i,a]\!\leftarrow\!\mathbf{1}[i<n]$triangleright draw $n$, initialize masks
2:$C_{1}\!\sim\!\mathcal{N}(0,I)$;  $\tilde{C}_{1}\!\leftarrow\!C_{1}-\bar{C}_{1,S_{1}}$triangleright center Gaussian prior by initial mask
3:for $t=1$ down to 00 do
4:     $\tilde{C}_{t}\!\leftarrow\!C_{t}-\bar{C}_{t,S_{t}}$;
6:     $\tilde{L}^{E}_{t}\!\leftarrow\!\textsc{SampleEdges}(L^{E}_{t},n)$triangleright enforce one parent per building block (Algorithm 2)
7:     $X_{t-\Delta t}\!\leftarrow\!\textsc{CatSample}(L^{X}_{t})$; $E_{t-\Delta t}\!\leftarrow\!\textsc{CatSample}(\tilde{L}^{E}_{t})$triangleright take reverse step (Section B.6)
10:end for
11:$(L^{X},L^{E},\hat{\tilde{C}}_{0})\!\leftarrow\!f_{\theta}(X_{0},E_{0},\tilde{C}_{0},n,0)$triangleright final deterministic denoise ($t=0$)
12:$\hat{X}_{0}\!\leftarrow\!\arg\max_{k}L_{\theta}^{X}[\cdots,k];\,$ $\hat{E}_{0}\!\leftarrow\!\arg\max_{k}L_{\theta}^{E}[\cdots,k];\,$ $\hat{C}_{0}\!\leftarrow\!\hat{\tilde{C}}_{0}-\bar{\hat{\tilde{C}}}_{0,S_{0}}$
13:return $(\hat{X}_{0},\hat{E}_{0},\hat{C}_{0})$
B.5 Building Block Logit Predictions
The SemlaFlow[Irwin et al., 2025] backbone outputs atom–atom edge features $E^{\mathrm{atom}}_{\theta}\in\mathbb{R}^{B\times(NM)\times(NM)\times d_{\text{edge}}}$. To obtain building block-level tensors, we apply two parallel 2-D convolutions (one for nodes, one for edges) with stride $M$, followed by MLP classifiers that map the pooled features back to their original one-hot vocabularies.
Stride-pooled convolution.
[…]
Node head.
We flatten $X_{\text{pool}}$ along its channel axis, concatenate the residual building block one-hot matrix $X_{t}$, and pass the result through a two-layer MLP to obtain
$L_{\theta}^{X_{t}}\in\mathbb{R}^{B\times N\times|\mathcal{B}|}.$ (13)
Edge head.
We concatenate $E_{\text{pool}}$ with the residual building block-edge one-hot tensor $E_{t}$, apply an analogous two-layer MLP, and symmetrize to produce
Atom Features.
The SemlaFlow[Irwin et al., 2025] backbone additionally outputs atom-level node features $X^{\mathrm{atom}}_{\theta}\in\mathbb{R}^{B\times(NM)\times d_{\text{node}}}$, which are incorporated into $E^{\mathrm{atom}}_{\theta}$ via a bond refinement message-passing layer. We find that extracting both building block and edge logits directly from the refined features $E^{\mathrm{atom}}_{\theta}$ marginally improves performance relative to separately predicting $L_{\theta}^{X_{t}}$ from $X^{\mathrm{atom}}_{\theta}$ and $L_{\theta}^{E_{t}}$ from $E^{\mathrm{atom}}_{\theta}$.
B.6 Discrete Noising Scheme
Following [Sahoo et al., 2024], we adopt an absorbing (masked) state noising scheme for $X_{0}$ and $E_{0}$:
where $(\alpha_{t})_{t\in[0,1]}$ is the monotonically decreasing noise schedule introduced in Section 2.
Reverse categorical posterior.
For node identities, we have
and, analogously, for edge labels
$q\!\left(E_{s}\mid E_{t},\,E_{0}\right)=\begin{cases}\operatorname{Cat}(E_{s};E_{t}),&E_{t}\neq\ \pi_{E}\\[6.0pt] \displaystyle\operatorname{Cat}\bigl{(}E_{s};\frac{(1-\alpha_{s})\pi_{E}+\alpha_{s}\,E_{\theta}^{t}}{1-\alpha_{t}}\bigr{)},&E_{t}=\pi_{E},\\[12.0pt] \end{cases}\!$ (17)
where $s<t$. Equations 16 and 17 are the direct translation of the reverse denoising process described by [Sahoo et al., 2024] into SynCoGen’s node–edge representation.
B.7 Noise Schedule Parameterization
Following MDLM [Sahoo et al., 2024], we parameterize the discrete noising schedule via $\alpha_{t}=e^{-\sigma(t)}$, where $\sigma(t):[0,1]\to\mathbb{R}^{+}$. In all experiments, we adopt the linear schedule:
$\sigma(t)=\sigma_{\max}t,$ (18)
Edge Symmetrization.
After drawing the upper-triangle entries of the one-hot edge tensor $E_{s}$ in either the forward or reverse (de)noising process, we enforce symmetry by copying them to the lower triangle:
$E_{s,jie}\;=\;E_{s,ije},\qquad 0\leq i<j<n,\;\;e\in\mathcal{R}V_{\max}^{2}.$
B.8 Positional Embeddings
Though SemlaFlow[Irwin et al., 2025] is permutationally invariant by design with respect to atom positions, SynCoGen dataset molecules require that atom order be fixed and grouped by building block for reconstruction purposes. To enforce this during training, we intentionally break permutation invariance by generating and concatenating to each input coordinate sinusoidal positional embeddings representing both global atom index and building block index.
B.9 Hyperparameters
We train SynCoGen for 100 epochs with a batch size of 128 and a global batch size of 512. Note that SemlaFlow and Midi are trained for 200 epochs, and EQGAT-diff is trained for up to 800 epochs. All models are trained with a linear noise schedule (see Section B.7), with the SUBS parameterization enabled. During training, a random conformer for each molecule is selected, then centered and randomly rotated to serve as the ground-truth coordinates $C_{0}$. All atomic coordinates are normalized by a constant $Z_{c}$ describing the standard deviation across all training examples. For the pairwise distance loss $\mathcal{L}_{pair}$, we set $d$ to 3Å, adjusted for normalization. During training, for each recentered input-prior pair $(\tilde{C}_{1},\tilde{C}_{0})$ we rotationally align $C_{1}$ to $C_{0}$. When training with noise scaling and the bond loss time threshold, we set the noise scaling coefficient to 0.2 and the time threshold to 0.25, above which bond length losses are zeroed. When training with auxiliary losses, we set the weights for the pairwise, sLDDT, and bond length loss components to 0.4, 0.4, and 0.2, respectively.
B.10 Computational Resources Used
We train all models on 2 H100-80GB GPUs.
B.11 Training Losses
Here, we define several loss terms that have proved useful for stabilizing training on 3-D geometry.
Graph loss.
Let $X_{0}$ and $E_{0}$ be the clean node and edge tensors. Following the MDLM implementation [Sahoo et al., 2024], we weigh the negative log-likelihood at step $t$ by
i.e. the cross-entropy between the one-hot ground truth and the predicted distributions for both nodes and edges.
MSE loss.
Pairwise loss.
where $d$ is the distance cut-off for pairwise terms. The default total loss value for the model is therefore
Smooth-LDDT loss [Abramson et al., 2024a].
Let $d_{ij}^{0}:=\lVert C_{0}[i]-C_{0}[j]\rVert_{2}$ and $d_{ij}^{\text{pred}}:=\lVert\hat{C}_{0}[i]-\hat{C}_{0}[j]\rVert_{2}$ be ground-truth and predicted inter-atomic distances, respectively. For each pair of atoms within a $15$ cutoff in the reference structure, we compute the per-pair score
where $\sigma(x)=1/(1+e^{-x})$ is the logistic function. The smooth-LDDT loss averages $1-\operatorname{sLDDT}_{ij}$ over all valid pairs,
Bond-length loss.
Given a set of intra-fragment bonds $\mathrm{bonds}=\{(p,q)\}$ extracted from the vocabulary, we penalize deviations in predicted bond lengths:
Self-Conditioning.
The modified SemlaFlow [Irwin et al., 2025] backbone operates on node and edges features at the atomic level, but outputs unnormalized prediction logits $\hat{X_{0}}\in\{0,1\}^{N\times|\mathcal{B}|}$ and $\hat{E_{0}}\in\;\{0,1\}^{N\times N\times|\mathcal{R}|V_{max}^{2}}$. We therefore implement modified self-conditioning for SynCoGen that projects previous step graph predictions $\hat{X_{0}}_{cond}$ and $\hat{E_{0}}_{cond}$ to the shape of $X^{atom}_{t}$ and $E^{atom}_{t}$ using an MLP.
B.12 Conformer generation
We randomly assembled 50 molecules with the reaction graph and used the standard conformational search (iMTD-GC) in CREST with GFN-FF to find all reference conformers. For both SynCoGen and RDKit ETKDG, we sampled 50 conformers per molecule and computed the coverage and matching scores. We used a relatively strict RMSD threshold of $\tau=0.75\ \mathrm{\AA}$.
Formally, COV is defined as:
B.13 Molecular inpainting
For the inpainting experiments in Section 5.2, we keep two fragments $\mathcal{D}=\{\mathcal{D}^{(1)},\mathcal{D}^{(2)}\}$ and their coordinates fixed and sample the remaining part of the molecule. We follow Section B.4 and initialize the graph prior $X_{1}$ with the one-hot encoding of the desired fragment $i$ at a specified node index in the graph (decided at random or based on the structure of the original molecule, so that it matches its scaffold). For each denoised fragment $\mathcal{D}^{(i)}$, we replace its coordinates at each time $t>0.03$ during sampling by
$C_{t}^{(i)}\;=\;(1-t)\,\tilde{C}_{0}^{(i)}\;+\;t\,\tilde{C}_{1}^{(i)},$
where $\tilde{C}_{0}^{(i)}$ and $\tilde{C}_{1}^{(i)}$ are the centered ground-truth and prior coordinates of fragment $i$, respectively, and all other fragments are updated as shown in Section B.4. For any $t\leq 0.03$, which for 100 sampling steps amounts to the last three steps in the path, we follow normal Euler steps as shown in Section B.4 to allow a refinement of the fixed coordinates in line with the rest of the predicted ones for the rest of the fragments. We empirically observed that this led to molecules with lower average energies.
Appendix C Baseline comparisons.
SemlaFlow
We evaluated SemlaFlow using the sampling script in the official codebase on GitHub [^1] https://github.com/rssrwn/semla-flow/, available under the MIT License. We reported results for a model trained on the GEOM [Axelrod and Gomez-Bombarelli, 2022] dataset (by sampling from the checkpoints provided in the repository) and from a model trained on our dataset (see Table 1). We trained SemlaFlow using the default hyperparameters for 150 epochs on a single conformer per molecule.
EQGAT-diff, MiDi, JODO, FlowMol
Appendix D Extended results and discussion
D.1 Sampling Ablations
We additionally follow FoldFlow to implement inference annealing, a time-dependent scaling on Euler step size that was found to empirically improve in-silico results in protein design Bose et al. [2024]. We studied multiplying the Euler step size at time $t$ by $5t$, $10t$, and $50t$. In practice, we employ $10t$ for our experiments unless otherwise noted.
Table 3: Sampling ablations. Results are averaged over 1000 generated samples, except retrosynthesis solve rate (out of 100). All results shown are at 50 epochs rather than 100 epochs in Table 1.
[…]
D.2 Metrics
@@
For synthesizability evaluation, we used the public AiZynthFinder and Syntheseus models. Due to the speed of these models, we only evaluate 100 randomly sampled generated examples. For AiZynthFinder, we used the USPTO policy, the Zinc stock, and we extended the search time to 800 seconds with an iteration limit of 200 seconds. For Syntheseus, we used the LocalRetro model with Retro* search under default settings, with Enamine REAL strict fragments as the stock. We additionally appended our building blocks as the stock, but found no meaningful difference in solved rates, presumably as most of our building blocks are already in the utilized stock. We note that we replaced all boranes with boronic acids due to simplifications made in our modeling (see Section A.2).
For energy evaluation, all results are from single-point calculations. For GFN-FF, we report the total energy minus the bond energies (equivalent to the sum of angle, dihedral, bond repulsion, electrostatic, dispersion, hydrogen bond, and halogen bond energies) as the intramolecular non-bond energies, and average it over the number of atoms. For GFN2-xTB, we report the dispersion interaction energies as the intramolecular non-covalent energies. We note that the total energies and bonded energies follow very similar trends. We note that MMFF94 energies are not parameterized for boron; therefore, we report them only for the Wasserstein distances in Table 4 and inpainting task in Table 6. Figures 3 and 7 show distributions obtained from 1,000 molecules generated by each generative method, along with 50,000 subsampled molecules from their respective training datasets. Gaussian kernel density estimation (bandwidth = 0.15) was used for linear distributions, while von Mises kernel density estimation ($\kappa=25$) was applied for circular distributions. Wasserstein-1 distances (computed linearly for lengths and energies, and on the circle for angles and dihedrals) were calculated using the Python Optimal Transport Package [Flamary et al., 2021].
D.3 De novo 3D molecule generation
Additional sample molecules and their synthetic pathways can be found in Figures 8 and 11. We further take a subset of random molecules and perform GFN2-xTB geometry optimization, and find that our generated samples closely recapitulate the optimized geometries (Figure 9). We also report the Wasserstein-1 distances and Jensen-Shannon divergence for the generative models on bond lengths, angles, and energies (Table 4), and show additional statistics in Figure 7. Lastly, we show the model’s ability to generate conformers with given graphs instead of co-generation in  Table 5.
Figure 7: Additional conformer bond length, angle, and dihedral distribution comparisons. a-b) Bond lengths, c-d) bond angles, e-f) dihedral angles. Solid curves denote training data densities; lower subpanels in (a-d) show deviations between generated samples and data.
@@ D.3 De novo 3D molecule generation
D.4 Molecular inpainting experiments
@@
In addition to the experiments in Section 5.2, we evaluate SynCoGen’s conditional sampling performance for the fragment linking framework against the state-of-the-art model DiffLinker [Igashov et al., 2024]. While DiffLinker is trained for fragment-linking, our model performs zero-shot fragment linking without any finetuning. For both models, the size of the linker was chosen so that it matches that of the original ligand: 2 extra nodes were sampled for SynCoGen and 15 linking atoms for DiffLinker in the case of 5L2S, while 3 extra nodes and 25 linking atoms were sampled for 4EYR and 7N7X. We specified leaving groups (for SynCoGen) and anchor points (for DiffLinker) so that the fragments are linked at the same positions as in the ligand. Results are shown in Table 6. No retrosynthetic pathways were found for the molecules in DiffLinkers, while SynCoGen models synthetic pathways and synthetic pathways can be easily drawn, with examples for 4EYR shown in Figure 11. This out-of-distribution task for SynCoGen leads to fewer valid molecules; however, for the valid candidates, SynCoGen has lower interaction energies and achieves 100% connectivity as it uses reaction-based assembly, whereas DiffLinker can sample disconnected fragments.
Table 6: Molecular inpainting task. Results are averaged over 1000 generated samples, except retrosynthesis solve rate (out of 100).
Method Target AiZyn. $\uparrow$ Synth. $\uparrow$ Valid. $\uparrow$ Connect. $\uparrow$ MMFF $\downarrow$ GFN-FF $\downarrow$ GFN2-xTB $\downarrow$ Diversity $\uparrow$ PB $\uparrow$

### v2

@@ D.2 Sampling Ablations
To examine the effect of de-noising edges probabilistically ("Default") against exclusively selecting the highest-probability edge ("Argmax") during sampling, we sample 1000 molecules unconditionally for each setting.
Table 5: Comparison of sampling strategies. All results are at 100 epochs (same with Table˜1) but without inference annealing.
Sampling Mode Validity $\uparrow$ PB $\uparrow$ Diversity $\uparrow$ Novelty (%) $\uparrow$
Default 0.947 0.8425 0.7811 95.24
Argmax 0.956 0.8483 0.7797 93.41
We see there is relatively insignificant differences (small improvement in the proportion of valid molecules is slightly higher in the "Argmax" setting, at the cost of slightly lower diversity and novelty). Note that the proportion of molecules with 3, 4 and 5 building blocks are sampled according to their respective distributions in SynSpace.
D.3 Larger Vocabulary
Section˜D.3 show unconditional generation results after training SynCoGen on SynSpace-Lfor 50 epochs, and as compared to SynCoGen on SynSpace for 50 epochs. We note that given a thousand fold increase in search space, longer training would typically be expected, and may further improve the results. Under this conservative setting, scaling to the larger search space yields very similar overall behavior. We observe a small decrease in RDKit validity and a moderate decrease in PoseBusters validity, and the molecules still retain good conformer energies and high retrosynthesis solve rates. These indicate the molecules are synthesizable, and the local geometry remains reasonable. Diversity is essentially unchanged, while novelty increases substantially from 94.2% to 99.1%, showing that the enlarged vocabulary is effectively used to explore new regions of chemical space.
@@ D.3 Larger Vocabulary
D.4 Metrics
@@
For synthesizability evaluation, we used the public AiZynthFinder and Syntheseus models. Due to the speed of these models, we only evaluate 100 randomly sampled generated examples. For AiZynthFinder, we used the USPTO policy, the Zinc stock, and we extended the search time to 800 seconds with an iteration limit of 200 seconds. For Syntheseus, we used the LocalRetro model with Retro* search under default settings, with Enamine REAL strict fragments as the stock. We additionally appended our building blocks as the stock, but found no meaningful difference in solved rates, presumably as most of our building blocks are already in the utilized stock. We note that we replaced all boranes with boronic acids due to simplifications made in our modeling (see Section˜A.2).
For energy evaluation, all results are from single-point calculations. For GFN-FF, we report the total energy minus the bond energies (equivalent to the sum of angle, dihedral, bond repulsion, electrostatic, dispersion, hydrogen bond, and halogen bond energies) as the intramolecular non-bond energies, and average it over the number of atoms. For GFN2-xTB, we report the dispersion interaction energies as the intramolecular non-covalent energies. We note that the total energies and bonded energies follow very similar trends. We note that MMFF94 energies are not parameterized for boron; therefore, we report them only for the Wasserstein distances in Table˜6 and inpainting task in Table˜8. Figures˜3 and 14 show distributions obtained from 1,000 molecules generated by each generative method, along with 50,000 subsampled molecules from their respective training datasets. Gaussian kernel density estimation (bandwidth = 0.15) was used for linear distributions, while von Mises kernel density estimation ($\kappa=25$) was applied for circular distributions. Wasserstein-1 distances (computed linearly for lengths and energies, and on the circle for angles and dihedrals) were calculated using the Python Optimal Transport Package (Flamary et al., 2021).
D.5 De novo 3D molecule generation
Figure 14: Additional conformer bond length, angle, dihedral, and energy distribution comparisons. a-b) Bond lengths, c) GFN2-xTB energy distribution, d-f) bond angles, g-h) dihedral angles. Solid curves denote training data densities; lower subpanels show deviations between generated samples and data.
@@ D.5 De novo 3D molecule generation
D.6 Molecular inpainting experiments
@@
In addition to the experiments in Section˜5.2, we evaluate SynCoGen’s conditional sampling performance for the fragment linking framework against the state-of-the-art model DiffLinker (Igashov et al., 2024). While DiffLinker is trained for fragment-linking, our model performs zero-shot fragment linking without any finetuning. For both models, the size of the linker was chosen so that it matches that of the original ligand: 2 extra nodes were sampled for SynCoGen and 15 linking atoms for DiffLinker in the case of 5L2S, while 3 extra nodes and 25 linking atoms were sampled for 4EYR and 7N7X. We specified leaving groups (for SynCoGen) and anchor points (for DiffLinker) so that the fragments are linked at the same positions as in the ligand. Results are shown in Table˜8. No retrosynthetic pathways were found for the molecules in DiffLinker, while SynCoGen models synthetic pathways and synthetic pathways can be easily drawn, with examples for 4EYR shown in Figure˜18. This out-of-distribution task for SynCoGen leads to fewer valid molecules; however, for the valid candidates, SynCoGen has lower interaction energies and achieves 100% connectivity as it uses reaction-based assembly, whereas DiffLinker can sample disconnected fragments.
Table 8: Molecular inpainting task. Results are averaged over 1000 generated samples, except retrosynthesis solve rate (out of 100). SynCoGen-FT denotes a light fine-tuning model for 5 epochs on in-painting with randomly fixed fragments from SynSpace.
Method Target AiZyn. $\uparrow$ Synth. $\uparrow$ Valid. $\uparrow$ Connect. $\uparrow$ MMFF $\downarrow$ GFN-$F$ $\downarrow$ GFN2-xTB $\downarrow$ Diversity $\uparrow$ PB $\uparrow$
@@
SynCoGen-FT 5L2S 77 84 75.3 100 4.25 6.58 -0.81 0.632 56.2
4EYR 42 78 62.0 100 10.13 5.33 -0.78 0.604 19.8
7N7X 57 77 73.6 100 4.09 6.86 -0.83 0.664 47.9
Table 9: Percentage of hard-to-synthesize chemical features in generated “valid” molecules from SynCoGen versus DiffLinker in fragment linking (out of 1000). Exotic bonds include hydrazine, nitro, nitramine, azide, diazo, peroxide, nitrate ester, fulminate. Fused large/small rings are where a fused ring contains a sub-ring that is larger than 6 atoms or smaller than 5 atoms.
DiffLinker SynCoGen
Chemical features 5L2S 3NDX 7N7X 5L2S 3NDX 7N7X
Macrocycles (>=9) 1.0% 72.6% 12.6% 0.0% 0.0% 0.0%
Fused rings with large/small rings 13.3% 81.4% 37.0% 0.0% 0.0% 0.0%
Large rings (7,8) 12.1% 9.9% 22.2% 0.0% 0.0% 0.0%
Disconnected 4.9% 18.1% 25.4% 0.0% 0.0% 0.0%
Exotic bonds 0.2% 1.2% 1.7% 0.0% 0.0% 0.0%
Total problematic % 22.8% 86.0% 61.5% 0.0% 0.0% 0.0%
@@
D.7 Pharmacophore-conditioned generation experiments
Table 10: Percentage of hard-to-synthesize chemical features in pharmacophore generation for CGFlow-ZS, Shepherd, Synformer, and SynCoGen 100 per target, 10 targets in total). Exotic bonds include hydrazine, nitro, nitramine, azide, diazo, peroxide, nitrate ester, fulminate. Fused large/small rings are where a fused ring contains a sub-ring that is larger than 6 atoms or smaller than 5 atoms.
Chemical features CGFlow-ZS Shepherd Synformer SynCoGen
Macrocycles (>=9) 0.0% 4.7% 1.8% 0.0%
Fused rings with large/small rings 31.1% 39.2% 1.9% 0.1%
Large rings (7,8) 1.2% 31.2% 4.0% 0.0%
Disconnected 0.0% 0.0% 0.0% 0.0%
Exotic bonds 0.0% 0.3% 1.3% 0.0%
Total problematic % 31.3% 46.8% 8.3% 0.1%
Figure 19: Docking score box-plot comparisons on pharmacophore-conditioned SynCoGen samples, randomly selected SynSpace samples, and randomly selected FDA-approved small molecules (100 for each target). Pharmacophore-conditioned SynCoGen  outperforms  SynSpace, which outperforms FDA-approved molecules. These results suggest that the reason why pharmacophore-conditioned SynCoGen can outperform other baselines may partially stem from the careful curation of building blocks, as SynSpace samples perform well in docking experiments. Lastly, we caution that docking is a merely a proxy for binding affinity, and we emphasize that the primary results are that SynCoGen  generates synthesizable molecules with reasonable poses when conditioned on pharmacophore profiles. Note all SynCoGen sampling runs were performed using a building block count fixed to 3.
Figure 20: Pharmacophore-conditioning task. Examples of docked SynCoGen-generated molecules (green) overlaid with PDB ligands (magenta) in their crystal structure pose.